OpenStudy (anonymous):

Find the perimeter of a rectangle with l = 3.2 m and w = 0.5 m. A. 1.6 m B. 3.2 m C. 3.7 m D. 7.4 m

OpenStudy (anonymous):

i know this is stupid but i need help i will fan and medal

OpenStudy (anonymous):

Perimeter = l + l + w + w = 3.2 + 3.2 + 0.5 + 0.5 = 3.2 x 2 + 0.5 x 2 = 6.4 + 1 =??

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

7.4 THANKS GUYS

OpenStudy (anonymous):

Np